An object travelling in a straight line away from a motion sensor at a constant rate moves from 4 m away to 16 m away between t = 1 s and t = 6 s. Describe the segment corresponding to this walk on the speed versus time graph.
Answer by mananth(16946)

An object travelling in a straight line away from a motion sensor at a constant rate moves
from 4 m away to 16 m away between t = 1 s and t = 6 s.
Speed = (Change in Displacement ) / (Change in Time)
= (16-4)/(6-1)= 12/5
the object is moving away at a constant speed of 2.4 meters per second
Now if you draw a graph of speed Vs Time, graph will be a horizontal line.
This line represents the object as it moves from 4 meters to 16 meters away from the sensor.